# Evaluating small vision-language models as AI assistants for radio astronomical source analysis tasks

### v2

@@ Abstract
This study explores small-scale Vision-Language Models (VLMs) as AI assistants for radio astronomy, combining LLM capabilities with vision transformers. We fine-tuned the LLaVA VLM on a dataset of 59k radio images from multiple surveys, enriched with 38k image-caption pairs from the literature. The fine-tuned models show clear improvements over base models in radio-specific tasks, achieving $\sim$30F1-score gains in extended source detection, but they underperform pure vision models and exhibit $\sim$20drop on general multimodal tasks. Inclusion of caption data and LoRA fine-tuning enhances instruction-following and helps recover $\sim$10accuracy on standard benchmarks.
@@ 1 Introduction
The upcoming Square Kilometer Array (SKA) (Dewdney et al., 2016) and its precursor telescopes (e.g., MeerKAT, ASKAP, LOFAR) are revolutionizing radio astronomy, enabling to probe the radio sky at unprecedented sensitivities and angular resolutions. SKA, once operational, is expected to produce exabytes of data annually. The immense volume and complexity of the generated data will challenge traditional data-processing methods, necessitating advanced computational and AI techniques to automate repetitive, resource-intensive tasks. In this context, deep-learning methodologies have already shown promising results in various analysis tasks including: source detection in 2D radio maps (Mostert et al., 2022; Zhang et al., 2022; Yu et al., 2022; Riggi et al., 2023; Lao et al., 2023; Cornu et al., 2024; Stuardi et al., 2024), source and host galaxy detection from 2D radio+IR maps (Wu et al., 2019; Gupta et al, 2023) or radio+optical maps (Lou et al., 2023), source detection in HI cubes (Liang et al., 2023; Håkansson et al., 2023; Barkai et al., 2023), source classification (Aniyan & Thorat, 2017; Tang et al., 2019; Ma et al., 2019; Maslej-Krešňáková et al., 2021; Tang et al., 2022; Nair et al., 2022; Riggi et al., 2024a), search for objects with peculiar morphology (Ralph et al., 2019; Galvin et al., 2020; Mostert et al., 2021; Gupta et al, 2022; Mesarcik et al., 2023; Lochner et al., 2023; Riggi et al., 2024b), fast radio burst detection (Connor & van Leeuwen, 2018; Agarwal, 2020), radio imaging (Schmidt et al., 2022; Geyer et al., 2023; Chiche et al., 2023), synthetic data generation (Rustige et al., 2023; Sortino et al., 2024; Martínez et al., 2024), and many others. The full potential of developed models, especially those using supervised learning techniques, is often hampered by the scarcity of large and balanced annotated radio datasets. Additionally, existing radio models typically employ data labeling schemes that vary widely across different analysis cases, hindering the integration of individual datasets into larger collections and restricting model usability beyond their initial applications.
@@
Recent studies (Slijepcevic et al., 2024; Riggi et al., 2024b; Lastufka et al., 2024) have sought to address the challenges posed by limited annotated training datasets by applying self-supervised learning (SSL) techniques, which utilize the extensive collections of unlabeled radio images available in current and past surveys. Several foundational radio models have been developed to effectively enable feature extraction from radio maps for a variety of tasks, such as data inspection, source extraction and classification, anomaly detection, and image retrieval. Pretrained SSL models have also been fine-tuned on smaller annotated datasets and specialized for these applications. Ongoing research is focusing on various areas: comparing alternative SSL methods on radio data (Cecconello et al., 2024), assessing the performance of SSL models pretrained on non-radio data (such as natural or optical images) for radio-specific tasks and vice versa (Riggi et al., 2024c; Lastufka et al., 2024b), exploring optimized dataset curation strategies, scaling up model training to larger architectures and millions of radio images, and defining more constraining downstream datasets and tasks.
@@
Large Language Models (LLMs) like GPT-4 (OpenAI, 2023), Claude3 (Anthropic, 2024), and open-source alternatives such as LLaMA (Touvron et al., 2023) or InternLM (Cai et al., 2024) have proven effective as AI assistants, showing remarkable zero-shot learning capabilities when prompted unseen data or tasks across a wide array of fields, including astronomy (Tanoglidis & Jain, 2024). Specializing and deploying very large open-source models is, however, currently prohibitive in terms of the required computing resources (high GPU requirements, memory demands, and power consumption). Furthermore, in addition to textual interaction, there is an increasing demand for models capable of processing visual data, facilitating multimodal reasoning for tasks like analyzing complex astronomical images. Some initiatives, such as AstroLLaMA (Nguyen et al., 2023; Perkowski et al., 2024), have started to address this by developing astronomy-specific mid-size LLMs, though these efforts are still limited to text-based inputs. Commercial solutions add additional cost concerns, particularly in inference and fine-tuning expenses, as demonstrated by Sun et al. (2024) in the context of interpreting multi-band galaxy observations. Furthermore, while large models are well-suited for tasks requiring extensive general knowledge, this scope may be more than what is needed in astronomy, where specialized knowledge is essential. Given these considerations, current research has increasingly turned to adapting smaller LLMs (i.e., those with fewer than 10 billion parameters) for specific domains, as well as investigating multi-modal models capable of processing combined data inputs, such as text, images, and videos.
@@
In this context, we aim to explore recent, state-of-the-art, small-scale vision-language models (VLMs) to develop AI assistants tailored to radio astronomy. These models combine both visual and textual comprehension by integrating LLM capabilities with vision transformers (Dosovitskiy et al., 2021) for image processing. Typically, VLMs comprise two main components: a vision encoder transformer that extracts features from input images, and an LLM that generates textual responses from combined visual and textual input representations. VLMs offer promising solutions to the outlined challenges in two ways. First, they can manage tasks through a text-based interface, allowing astronomers to specify task details and expected response formats in natural language. This flexibility enables the use of contextual image-text examples, making VLMs more adaptable than traditional vision models that are limited to predefined label outputs for specific tasks. Second, VLMs potentially support the integration of specialized vision encoder models that have been trained on unlabeled radio data through self-supervised learning (SSL) methods. This approach bridges the gap with ongoing SSL research, facilitating the reuse of existing radio SSL models and enabling the full utilization of extensive unlabeled image datasets from recent and past radio surveys. These smaller models offer a promising alternative, with lower computational costs and more manageable deployment requirements. Moreover, their specialized capabilities could be more than sufficient for the specific needs of radio astronomy, without the added complexity and resource demands of larger models. However, the suitability of these smaller models for astronomical tasks remains largely unexplored, particularly in radio source analysis.
@@
This paper seeks to evaluate the current state of small multimodal language models as AI assistants for radio astronomy. By investigating their strengths, limitations, and applicability to radio astronomical source analysis tasks, we hope to familiarize the community with this emerging technology and its potential, as well as to highlight the challenges that need to be addressed in future developments. A tailored VLM for radio astronomy could eventually assist astronomers in efficiently analyzing radio images without requiring extensive technical expertise in AI models. By leveraging a text-based interface, astronomers can perform complex and diverse image analysis tasks, even guiding the model using image-based examples. Key applications include the automated identification and retrieval of specific classes of radio sources in survey image data, as well as data quality assessment - enhancing the efficiency and accessibility of radio survey analysis. For instance, Gupta et al (2025) introduced a tool, EMUSE, built on a fine-tuned OpenCLIP [^1] https://github.com/mlfoundations/open_clip model, which enables users to search EMU survey data using either textual queries or template image similarity. Additionally, a VLM-based assistant could be deployed to support less experienced users (e.g., students, citizens) in ongoing crowdsourcing projects like the EMU Radio Galaxy Zoo [^2] https://www.zooniverse.org/projects/hongming-tang/radio-galaxy-zoo-emu. By providing real-time guidance and explanations, the model could help users classify radio sources, identify peculiar objects, and improve the reliability of crowdsourced annotations.
@@
The paper is organized as follows. Section 2 provides an overview of vision-language models, with a focus on the architecture of a prominent model, LLaVA, which we aim to adapt for radio-astronomical data. In Section 3, we describe our adapted model, termed radio-llava, including the training datasets and methodology. Section 4 presents the evaluation of the specialized model across several radio-astronomy tasks. Finally, Section 5 summarizes the results and discusses directions for future research.
@@ 2.1.1 Model overview
LLaVA (Large Language and Vision Assistant) (Liu et al., 2023) is a state-of-the-art multimodal model that integrates both visual and textual understanding, combining the capabilities of large language models (LLMs) with vision processing abilities. Its primary function is to interpret and generate responses to input that includes both images and text, making it ideal for tasks like visual question answering (VQA), image captioning, and other vision-language tasks. Since the first release, the model demonstrated exceptional multimodal conversational skills, often displaying behavior comparable to GPT-4V when tasked with interpreting novel images and following new instructions for the first time.
Following releases (LLaVA 1.5, Liu et al. 2024a) greatly enhanced model capabilities by integrating a larger set of academic-focused instructional data, achieving state-of-the-art results on numerous benchmarks while utilizing a highly data-efficient strategy. Recent advancements in the LLaVA series, including models like LLaVA-NeXT (Liu et al., 2024b) and LLaVA-OneVision (Li et al., 2024), have significantly broadened the scope of input modalities they can handle, supporting both single or multiple images as well as video content. These improvements were driven by three key innovations: the AnyRes technique for processing high-resolution images, the expansion of high-quality instruction datasets, and the integration of the most advanced open-source LLMs available at the time, further enhancing model capabilities across diverse tasks. Various variants or specialization of the first LLaVA models have been produced so far. For example, TinyLLaVA (Zhou et al., 2024; Jia et al., 2024) is a compact refactored variant of the original LLaVA 1.5 model, designed to enable easier inclusion of alternative light vision and LLM models, thus significantly reducing overall model size and resource requirements. LLaVA-Med (Li et al., 2023a) is a specialized variant of the LLaVA model designed to assist in medical image analysis and diagnostics by fine-tuning its multimodal capabilities on medical datasets such as X-rays, MRIs, and other healthcare-related visual data.
@@ 2.1.2 Model architecture
The LLaVA model, schematically represented in Fig. 1, consists of these components:
  • Vision Encoder: Processes image data using a pre-trained Vision Transformer (ViT) model with multiple transformer layers, such as CLIP (Radford et al., 2021) or SigLIP (Zhai et al., 2023). It extracts visual features from input images;
  • Language Model: Handles text processing, typically an LLM such as Qwen (Bai et al., 2023), capable of understanding, generating, and reasoning with natural language;
@@
    The input image is divided into patches and passed through the vision encoder, which extracts key features such as objects, colors, textures, and spatial relationships. The projector then converts these visual representations into language-compatible embeddings;
@@ 3.1.1 Q&A dataset
  • Fine-Grained Datasets: These datasets, typically used for training radio object detection and segmentation models like YOLO (Redmon et al., 2016) or Mask R-CNN (He et al, 2017), contain wide-field images (a few arcminutes in size) with region- or pixel-level annotations, including object positions (centers, bounding boxes, segmentation masks), classification labels, and confidence scores.
@@
It is important to note that classification schemes vary across datasets. Some provide astrophysical source-type labels (e.g., HII, SNR, GALAXY), while others focus on morphological classifications (e.g., FR-I vs. FR-II, COMPACT vs. EXTENDED vs. DIFFUSE). Before generating the conversational Q&A dataset, we aimed to standardize terminology whenever possible. In some cases (e.g., see Section A.2), we performed cross-matching and relabeling to augment the original datasets by adding additional labels to certain images or objects. However, variations in labeling schemes remain unavoidable due to the lack of annotation standards in the radio astronomy community. In this respect, our goal is to fine-tune an LLM model that is exposed to diverse classification schemes, making it potentially adaptable to different user domains.
@@
    Automated Q&A Generation Using a Pretrained VLM: We generated multiple Q&A interactions per image using a InternVL VLM model [^4] We used InternVL2_5-8B-MPO model version, available here: https://huggingface.co/OpenGVLab/InternVL2_5-8B-MPO, https://github.com/OpenGVLab/InternVL (Chen et al., 2024). The model was fed with the image, template caption, and structured prompts to ensure that the generated conversations remained faithful to the original image and annotation content, and included at least the following questions:

### v3

# radio-llava: Advancing Vision-Language Models for Radio Astronomical Source Analysis

@@ Abstract
This study explores small-scale Vision-Language Models (VLMs) as AI assistants for radio astronomy, combining LLM capabilities with vision transformers. We fine-tuned the LLaVA VLM on a dataset of 59k radio images from multiple surveys, enriched with 38k image-caption pairs from the literature. The fine-tuned models show clear improvements over base models in radio-specific tasks, achieving $\sim$30F1-score gains in extended source detection, but they underperform vision-only classifiers and exhibit $\sim$20drop on general multimodal tasks. Inclusion of caption data and LoRA fine-tuning enhances instruction-following and helps recover $\sim$10accuracy on multimodal benchmarks (e.g., ChartQA/DocVQA).
@@ 1 Introduction
The upcoming Square Kilometer Array (SKA) Dewdney et al. (2016) and its precursor telescopes (e.g., MeerKAT, ASKAP, LOFAR) are revolutionizing radio astronomy, enabling to probe the radio sky at unprecedented sensitivities and angular resolutions. SKA, once operational, is expected to produce exabytes of data annually. The immense volume and complexity of the generated data will challenge traditional data-processing methods, necessitating advanced computational and AI techniques to automate repetitive, resource-intensive tasks.
@@
In this context, deep-learning methodologies have already shown promising results in various analysis tasks including: source detection in 2D radio maps Mostert et al. (2022); Zhang et al. (2022); Yu et al. (2022); Riggi et al. (2023); Lao et al. (2023); Cornu et al. (2024); Stuardi et al. (2024), source and host galaxy detection from 2D radio+IR maps Wu et al. (2019); Gupta et al (2023) or radio+optical maps Lou et al. (2023), source detection in HI cubes Liang et al. (2023); Håkansson et al. (2023); Barkai et al. (2023), source classification Aniyan & Thorat (2017); Tang et al. (2019); Ma et al. (2019); Maslej-Krešňáková et al. (2021); Tang et al. (2022); Nair et al. (2022); Riggi et al. (2024a), search for objects with peculiar morphology Ralph et al. (2019); Galvin et al. (2020); Mostert et al. (2021); Gupta et al (2022); Mesarcik et al. (2023); Lochner et al. (2023); Riggi et al. (2024b), fast radio burst detection Connor & van Leeuwen (2018); Agarwal (2020), radio imaging Schmidt et al. (2022); Geyer et al. (2023); Chiche et al. (2023), synthetic data generation Rustige et al. (2023); Sortino et al. (2024); Martínez et al. (2024), and many others. The full potential of developed models, especially those using supervised learning techniques, is often hampered by the scarcity of large and balanced annotated radio datasets. Additionally, existing radio models typically employ data labelling schemes that vary widely across different analysis cases, hindering the integration of individual datasets into larger collections and restricting model usability beyond their initial applications.
@@
Recent studies Slijepcevic et al. (2024); Riggi et al. (2024b); Lastufka et al. (2024) have sought to address the challenges posed by limited annotated training datasets by applying self-supervised learning (SSL) techniques, which utilize the extensive collections of unlabelled radio images available in current and past surveys. Several foundational radio models have been developed to effectively enable feature extraction from radio maps for a variety of tasks, such as data inspection, source extraction and classification, anomaly detection, and image retrieval. Pretrained SSL models have also been fine-tuned on smaller annotated datasets and specialized for these applications. Ongoing research is focusing on various areas: comparing alternative SSL methods on radio data Cecconello et al. (2024), assessing the performance of SSL models pretrained on non-radio data (such as natural or optical images) for radio-specific tasks and vice versa Riggi et al. (2024c); Lastufka et al. (2024b), exploring optimized dataset curation strategies, scaling up model training to larger architectures and millions of radio images, and defining more constraining downstream datasets and tasks.
@@
Large Language Models (LLMs) like GPT-4 OpenAI (2023), Claude3 Anthropic (2024), and open-source alternatives such as LLaMA Touvron et al. (2023) or InternLM Cai et al. (2024) have proven effective as AI assistants, showing remarkable zero-shot learning capabilities when prompted unseen data or tasks across a wide array of fields, including astronomy Tanoglidis & Jain (2024). Specializing and deploying very large open-source models is, however, currently prohibitive in terms of the required computing resources (high GPU requirements, memory demands, and power consumption). Furthermore, in addition to textual interaction, there is an increasing demand for models capable of processing visual data, facilitating multimodal reasoning for tasks like analysing complex astronomical images. Some initiatives, such as AstroLLaMA Nguyen et al. (2023); Perkowski et al. (2024), have started to address this by developing astronomy-specific mid-size LLMs, though these efforts are still limited to text-based inputs. Commercial solutions add additional cost concerns, particularly in inference and fine-tuning expenses, as demonstrated by Sun et al. (2024) in the context of interpreting multi-band galaxy observations. Furthermore, while large models are well-suited for tasks requiring extensive general knowledge, this scope may be more than what is needed in astronomy, where specialized knowledge is essential. Given these considerations, current research has increasingly turned to adapting smaller LLMs (i.e., those with fewer than 10 billion parameters) for specific domains, as well as investigating multi-modal models capable of processing combined data inputs, such as text, images, and videos.
@@
In this context, we aim to explore recent, state-of-the-art, small-scale vision-language models (VLMs) to develop AI assistants tailored to radio astronomy. These models combine both visual and textual comprehension by integrating LLM capabilities with vision transformers Dosovitskiy et al. (2021) for image processing. Typically, VLMs comprise two main components: a vision encoder transformer that extracts features from input images, and an LLM that generates textual responses from combined visual and textual input representations. VLMs offer promising solutions to the outlined challenges in two ways. First, they can manage tasks through a text-based interface, allowing astronomers to specify task details and expected response formats in natural language. This flexibility enables the use of contextual image-text examples, making VLMs more adaptable than traditional vision models that are limited to predefined label outputs for specific tasks. Second, VLMs potentially support the integration of specialized vision encoder models that have been trained on unlabelled radio data through self-supervised learning (SSL) methods. This approach bridges the gap with ongoing SSL research, facilitating the reuse of existing radio SSL models and enabling the full utilization of extensive unlabelled image datasets from recent and past radio surveys. These smaller models offer a promising alternative, with lower computational costs and more manageable deployment requirements. Moreover, their specialized capabilities could be more than sufficient for the specific needs of radio astronomy, without the added complexity and resource demands of larger models. However, the suitability of these smaller models for astronomical tasks remains largely unexplored, particularly in radio source analysis.
@@
This paper seeks to evaluate the current state of small multimodal language models as AI assistants for radio astronomy. By investigating their strengths, limitations, and applicability to radio astronomical source analysis tasks, we hope to familiarize the community with this emerging technology and its potential, as well as to highlight the challenges that need to be addressed in future developments. A tailored VLM for radio astronomy could eventually assist astronomers in efficiently analysing radio images without requiring extensive technical expertise in AI models. By leveraging a text-based interface, astronomers can perform complex and diverse image analysis tasks, even guiding the model using image-based examples. Key applications include the automated identification and retrieval of specific classes of radio sources in survey image data, as well as data quality assessment - enhancing the efficiency and accessibility of radio survey analysis. Additionally, a VLM-based assistant could be deployed to support less experienced users (e.g., students, citizens) in ongoing crowdsourcing projects like the EMU Radio Galaxy Zoo [^1] https://www.zooniverse.org/projects/hongming-tang/radio-galaxy-zoo-emu. By providing real-time guidance and explanations, the model could help users classify radio sources, identify peculiar objects, and improve the reliability of crowdsourced annotations.
Multi-modal models have only very recently begun to gain traction in astronomy - with most developments emerging within the past year - as a means to bridge heterogeneous data modalities such as images, spectra, and natural language. These efforts have primarily focused on adapting the CLIP (Contrastive Language–Image Pretraining) model Radford et al. (2021) and its derivatives to astronomical tasks, leveraging their ability to align visual and textual representations in a shared latent space. For instance, Gupta et al (2025) introduced EMUSE, a tool built on a fine-tuned OpenCLIP [^2] https://github.com/mlfoundations/open_clip model that enables users to search EMU survey data using either textual queries or template image similarity. A similar application is PAPERCLIP Mishra-Sharma et al. (2024), a CLIP-based model fine-tuned on Hubble Space Telescope (HST) proposal data -- including image observations and proposal abstracts -- which enables cross-modal retrieval based on image content or textual queries such as object names or scientific use cases. AstroCLIP Parker et al. (2024) presents a powerful CLIP-style foundation model for galaxies in the optical domain, where the textual modality in the original CLIP framework is replaced by galaxy spectra. The model aligns image and spectrum modalities in a shared embedding space after self-supervised pre-training of each modality independently, and demonstrates impressive performance on tasks such as morphology classification, redshift estimation, and property inference.
Unlike these studies, our approach focuses on instruction-following vision-language models like LLaVA (Large Language and Vision Assistant) Liu et al. (2023) to enable open-ended captioning, visual question answering, and multi-turn scientific dialogue grounded in domain-specific radio astronomy data. These tasks go beyond the static alignment capabilities of CLIP models, which lack generative, reasoning, and conversational abilities. Our model is thus particularly suited for exploratory analysis, educational interfaces, and assistant-style tools that can explain, summarise, or discuss diagnostic plots and observational data. Meanwhile, CLIP-based models remain better suited for scalable retrieval, zero-shot classification, and semantic similarity search over large datasets.
@@
The paper is organized as follows. Section 2 provides an overview of vision-language models, with a focus on the architecture of a prominent model, LLaVA, which we aim to adapt for radio-astronomical data. In Section 3, we describe our adapted model, termed radio-llava, including the training datasets and methodology. Section 4 presents the evaluation of the specialized model across several radio-astronomy tasks. Finally, Section 5 summarises the results and discusses directions for future research.
@@ 2.1.1 Model overview
LLaVA (Large Language and Vision Assistant) Liu et al. (2023) is a state-of-the-art multimodal model that integrates both visual and textual understanding, combining the capabilities of large language models (LLMs) with vision processing abilities. Its primary function is to interpret and generate responses to input that includes both images and text, making it ideal for tasks like visual question answering (VQA), image captioning, and other vision-language tasks. Since the first release, the model demonstrated exceptional multimodal conversational skills, often displaying behaviour comparable to GPT-4V when tasked with interpreting novel images and following new instructions for the first time.
Following releases (LLaVA 1.5, Liu et al. (2024a)) greatly enhanced model capabilities by integrating a larger set of academic-focused instructional data, achieving state-of-the-art results on numerous benchmarks while utilizing a highly data-efficient strategy. Recent advancements in the LLaVA series, including models like LLaVA-NeXT Liu et al. (2024b) and LLaVA-OneVision Li et al. (2024), have significantly broadened the scope of input modalities they can handle, supporting both single or multiple images as well as video content. These improvements were driven by three key innovations: the AnyRes technique for processing high-resolution images, the expansion of high-quality instruction datasets, and the integration of the most advanced open-source LLMs available at the time, further enhancing model capabilities across diverse tasks. Various variants or specialization of the first LLaVA models have been produced so far. For example, TinyLLaVA Zhou et al. (2024); Jia et al. (2024) is a compact refactored variant of the original LLaVA 1.5 model, designed to enable easier inclusion of alternative light vision and LLM models, thus significantly reducing overall model size and resource requirements. LLaVA-Med Li et al. (2023a) is a specialized variant of the LLaVA model designed to assist in medical image analysis and diagnostics by fine-tuning its multimodal capabilities on medical datasets such as X-rays, MRIs, and other healthcare-related visual data.
@@ 2.1.2 Model architecture
The LLaVA model, schematically represented in Figure 1, consists of these components:
  • Vision Encoder: Processes image data using a pre-trained Vision Transformer (ViT) model with multiple transformer layers, such as CLIP Radford et al. (2021) or SigLIP Zhai et al. (2023). It extracts visual features from input images;
  • Language Model: Handles text processing, typically an LLM such as Qwen Bai et al. (2023), capable of understanding, generating, and reasoning with natural language;
@@
    The input image is divided into patches and passed through the vision encoder, which extracts key features such as objects, colours, textures, and spatial relationships. The projector then converts these visual representations into language-compatible embeddings;
@@ 3.1.1 Q&A dataset
  • Fine-Grained Datasets: These datasets, typically used for training radio object detection and segmentation models like YOLO Redmon et al. (2016) or Mask R-CNN He et al (2017), contain wide-field images (a few arcminutes in size) with region- or pixel-level annotations, including object positions (centres, bounding boxes, segmentation masks), classification labels, and confidence scores.
@@
It is important to note that classification schemes vary across datasets. Some provide astrophysical source-type labels (e.g., HII, SNR, GALAXY), while others focus on morphological classifications (e.g., FR-I vs. FR-II, COMPACT vs. EXTENDED vs. DIFFUSE). Before generating the conversational Q&A dataset, we aimed to standardize terminology whenever possible. In some cases (e.g., see Section A.2), we performed cross-matching and relabelling to augment the original datasets by adding additional labels to certain images or objects. However, variations in labelling schemes remain unavoidable due to the lack of annotation standards in the radio astronomy community. In this respect, our goal is to fine-tune an LLM model that is exposed to diverse classification schemes, making it potentially adaptable to different user domains.
@@
    Automated Q&A Generation Using a Pretrained VLM: We generated multiple Q&A interactions per image using a InternVL VLM model [^4] We used InternVL2_5-8B-MPO model version, available here: https://huggingface.co/OpenGVLab/InternVL2_5-8B-MPO, https://github.com/OpenGVLab/InternVL Chen et al. (2024). The model was fed with the image, template caption, and structured prompts to ensure that the generated conversations remained faithful to the original image and annotation content, and included at least the following questions: